# Exploring black hole shadows in axisymmetric spacetimes with coordinate-independent methods and neural networks

### v1

@@ II.1 Ray Tracing
To derive the equations for the $t$- and $\phi$-components of the photon position, we exploit the circular spacetimes, which admit two standard Killing vectors: $\xi=(1,0,0,0)$ and $\eta=(0,0,0,1)$. These symmetries lead to the conservation of energy and angular momentum along the photon’s trajectory, expressed as:
@@
Each photon is traced backward in time from the observer’s image plane, located at a distance $d$ and inclined at an angle $i$ with respect to the black hole’s spin axis. The initial position and four-momentum in Boyer–Lindquist coordinates are then computed as Bambi (2017):
@@
$\left(\frac{dt}{d\lambda^{\prime}}\right)_{j}=-\frac{g_{t\phi}}{g_{tt}}\left(\frac{d\phi}{d\lambda^{\prime}}\right)_{j}+\sqrt{\frac{g_{t\phi}^{2}}{g_{tt}^{2}}\left(\frac{d\phi}{d\lambda^{\prime}}\right)_{j}^{2}-\left[\frac{g_{rr}}{g_{tt}}\left(\frac{dr}{d\lambda^{\prime}}\right)_{j}^{2}+\frac{g_{\theta\theta}}{g_{tt}}\left(\frac{d\theta}{d\lambda^{\prime}}\right)_{j}^{2}+\frac{g_{\phi\phi}}{g_{tt}}\left(\frac{d\phi}{d\lambda^{\prime}}\right)_{j}^{2}\right]},$ (13)
@@ II.2 Black Hole Shadow in Different Metrics and Observational Constraints
where $r_{\mathrm{sch}}=\sqrt{27}\;M$ is the radius of the shadow for a Schwarzschild black hole.
Figure 1 shows the behavior of $r_{\mathrm{sh}}$ and $\delta$ for the Kerr spacetime. As the spin parameter $a_{*}$ increases, the shadow areal radius $r_{\mathrm{sh}}$ gradually decreases. The maximum deviation from the Schwarzschild case occurs when the observer’s inclination angle is $0^{\circ}$ (the observer is along the black hole spin axis) and the black hole spin is near its maximum. Even in this extreme case, the deviation remains below approximately $7.5\%$ Takahashi (2004); Akiyama et al. (2022b). Similarly, Fig. 2 illustrates how variations in the parameter $\gamma$ of the $\gamma$–metric modify both the shadow radius and the Schwarzschild deviation. In this case, the deformation leads to deviations in $\delta$ ranging approximately from $-0.07$ to $0.06$.
@@
For the Taub-NUT metric (Fig. 5), observational data place an upper limit on the NUT parameter $l$: it must satisfy $0.440\leq l<$ 0.44 according to Keck and $0.20\leq l<$ 0.2 according to VLTI.
In the Kaluza-Klein metric (Fig. 6), considering the case where the magnetic and electric charges are equal ($b=c$), the allowed range for the parameter $b$ is $3.92\leq b<$ 3.9 for Keck and $4.422\leq b<$ 4.42 for VLTI observations.
[…]
@@
[…]
@@
[…]
@@ III Shadow Characterization: Coordinate-Independent Formalism
$R_{0}:=$ $\left(\int_{0}^{2\pi}R^{\prime}\mathrm{d}\psi^{\prime}\right)^{-1}\left[\left(\int_{0}^{2\pi}R^{\prime 2}\cos\psi^{\prime}\mathrm{d}\psi^{\prime}\right)^{2}\right.$ (16)
@@
$R:=$ ${\left[\left(R^{\prime}\cos\psi^{\prime}-R_{0}\cos\psi_{0}\right)^{2}\right.}$ (18)
@@
$\psi:=$ $\tan^{-1}\frac{R^{\prime}\sin\psi^{\prime}-R_{0}\sin\psi_{0}}{R^{\prime}\cos\psi^{\prime}-R_{0}\cos\psi_{0}}.$
@@
[…]
for the values of =0,, $\ell=0,\ ...,\ 5$.
@@
[…]
@@ A.5 Kaluza–Klein Metric
$\frac{H_{1}}{M^{2}}$ $=\frac{8(b-2)(c-2)b}{(b+c)^{3}}+\frac{4(b-2)x}{b+c}+x^{2}-\frac{2b\sqrt{(b^{2}-4)(c^{2}-4)}a_{*}\cos\theta}{(b+c)^{2}}+a_{*}^{2}\cos^{2}\theta,$ (47)
$\frac{H_{2}}{M^{2}}$ $=\frac{8(b-2)(c-2)c}{(b+c)^{3}}+\frac{4(c-2)x}{b+c}+x^{2}-\frac{2c\sqrt{(b^{2}-4)(c^{2}-4)}a_{*}\cos\theta}{(b+c)^{2}}+a_{*}^{2}\cos^{2}\theta,$

### v2

@@ II.1 Ray Tracing
To derive the equations for the $t$- and $\phi$-components of the photon position, we adopt a Boyer–Lindquist–like coordinate system $(t,r,\theta,\phi)$ with metric signature (,+,+,+)(-,+,+,+). In this convention, stationary and axisymmetric (circular) spacetimes admit two Killing vectors associated with time-translation and axial symmetry, respectively: $\xi^{\mu}=(1,0,0,0)$ and $\eta^{\mu}=(0,0,0,1)$. These symmetries lead to the conservation of energy and angular momentum along the photon trajectory, expressed as:
@@
Each photon is traced backward in time from the observer’s image plane (or “screen”), which we place at a large distance $d\gg M$ from the black hole (see Fig. 1). At such distances the spacetime curvature is negligible and the metric reduces to flat space, allowing the screen to be treated as a Cartesian plane. The screen is tilted by the inclination angle $i$ with respect to the black hole’s spin axis. We parametrize the screen using celestial coordinates $(\alpha,\beta)$, where $\alpha$ measures the displacement along the horizontal axis and $\beta$ along the vertical axis. These screen coordinates are directly related to the photon’s impact parameters and enter the initial conditions of the geodesics through Eqs. (7)–(13). The initial position and four-momentum in Boyer–Lindquist coordinates are then computed as Bambi (2017):
Figure 1: Ray-tracing setup: photons are launched from the observer’s screen at distance $d$, inclined by angle $i$ relative to the spin axis. The screen is parametrized by coordinates $(\alpha,\beta)$, which map to the photon impact parameters used in the geodesic integration.
@@
$\begin{split}&\left(\frac{dt}{d\lambda^{\prime}}\right)_{j}=-\frac{g_{t\phi}}{g_{tt}}\left(\frac{d\phi}{d\lambda^{\prime}}\right)_{j}\\ &\quad+\sqrt{\frac{g_{t\phi}^{2}}{g_{tt}^{2}}\left(\frac{d\phi}{d\lambda^{\prime}}\right)_{j}^{2}-\left[\frac{g_{rr}}{g_{tt}}\left(\frac{dr}{d\lambda^{\prime}}\right)_{j}^{2}+\frac{g_{\theta\theta}}{g_{tt}}\left(\frac{d\theta}{d\lambda^{\prime}}\right)_{j}^{2}+\frac{g_{\phi\phi}}{g_{tt}}\left(\frac{d\phi}{d\lambda^{\prime}}\right)_{j}^{2}\right]},\end{split}$
@@ II.2 Black Hole Shadow in Different Metrics and Observational Constraints
where $r_{\mathrm{sch}}=\sqrt{27}\,M$ is the shadow radius of a Schwarzschild black hole for an observer at infinity. In numerical ray-tracing we approximate this limit by placing the observer’s screen at a large but finite distance $d=10^{5}M$, which reproduces the observer-at-infinity result to high accuracy.
Figure 2 shows the behavior of $r_{\mathrm{sh}}$ and $\delta$ for the Kerr spacetime. As the spin parameter $a_{*}$ increases, the shadow areal radius $r_{\mathrm{sh}}$ gradually decreases. The maximum deviation from the Schwarzschild case occurs when the observer’s inclination angle is $0^{\circ}$ (the observer is along the black hole spin axis) and the black hole spin is near its maximum. Even in this extreme case, the deviation remains below approximately $7.5\%$ Takahashi (2004); Akiyama et al. (2022b). Similarly, Fig. 3 illustrates how variations in the parameter $\gamma$ of the $\gamma$–metric modify both the shadow radius and the Schwarzschild deviation. In this case, the deformation leads to deviations in $\delta$ ranging approximately from 0.07-0.07 to $0.06$.
@@
For the Taub-NUT metric (Fig. 6), observational data place an upper limit on the NUT parameter $l$: it must satisfy $0\leq l<0.44$ according to Keck and $0\leq l<0.2$ according to VLTI.
In the Kaluza-Klein metric (Fig. 7), considering the case where the magnetic and electric charges are equal ($b=c$), the allowed range for the parameter $b$ is $2\leq b<3.9$ for Keck and $2\leq b<4.42$ for VLTI observations.
Figure 2: Shadow areal radius $r_{\rm sh}$ (left) and Schwarzschild deviation parameter $\delta$ (right) of a Kerr black hole as functions of observer inclination angle $i$ for spin parameters $a_{*}=0.00,\,0.10,\,0.20,\,\dots,\,0.99$. Curves are colour-coded from blue ($a_{*}=0$) to dark red ($a_{*}=0.99$) and use alternating solid/dashed styles for clarity.
Figure 3: Same as Fig. 2 for the $\gamma$-metric ($\gamma=0.50$$1.50$).
@@
Figure 5: Same as Fig. 4 for the $\gamma$ metric at $i=0^{\circ},\,45^{\circ},\,90^{\circ}$.
@@
Figure 7: Same as Fig. 4 for the Kaluza-Klein metric: $r_{\rm sh}$ and $\delta$ vs. parameter $b$ for $a_{*}=0.0$$0.9$.
@@ III Shadow Characterization: Coordinate-Independent Formalism
$R_{0}$= $\left(\int_{0}^{2\pi}R^{\prime}\mathrm{d}\psi^{\prime}\right)^{-1}\left[\left(\int_{0}^{2\pi}R^{\prime 2}\cos\psi^{\prime}\mathrm{d}\psi^{\prime}\right)^{2}\right$. (16)
@@
Equation (16) defines the effective center $R_{0}$ of the shadow in a coordinate-independent way. The motivation for this construction is that the geometric center of the observed contour does not necessarily coincide with the observer’s coordinate origin $O^{\prime}$. Simply expanding $R^{\prime}(\psi^{\prime})$ around $O^{\prime}$ could therefore introduce spurious asymmetries. The integral expression for $R_{0}$ plays the role of a centroid, analogous to the center of mass of a closed curve, and provides a robust way to anchor the shadow’s effective center before carrying out further analysis.
@@
$R$= ${\left[\left(R^{\prime}\cos\psi^{\prime}-R_{0}\cos\psi_{0}\right)^{2}\right.}$ (18)
@@
$\psi$= $\tan^{-1}\frac{R^{\prime}\sin\psi^{\prime}-R_{0}\sin\psi_{0}}{R^{\prime}\cos\psi^{\prime}-R_{0}\cos\psi_{0}}.$
@@
$c_{\ell}=\frac{2\ell+1}{2}\int_{\lambda_{1}}^{\lambda_{2}}R(\lambda)P_{\ell}(\cos\psi)\sin\psi\left(\frac{d\psi}{d\lambda}\right)d\lambda\ ,$
for the values of $\ell=0,\ ...,\ 5$.
@@
In addition, Figure 10 presents heatmaps of the first four Fourier coefficients $C_{0}$$C_{3}$, plotted over the parameter space of spin $a_{*}$ and inclination angle $i$. The coefficient $C_{0}$ reflects the average size of the shadow and varies monotonically, while the higher-order coefficients encode features related to asymmetry and distortion. These patterns reveal how Fourier coefficients contain rich, structured information that can be used for inverse parameter estimation from observational shadow data.
Figure 9: Behavior of the shadow expansion magnitudes $c_{\ell}$ ($C_{\ell}$) illustrating their rapid decay. Top left: Legendre expansion magnitudes versus mode order $\ell$ for spins $a_{*}=0.4,0.6,0.8,0.95$. Top right: Legendre expansion magnitudes versus $a_{*}$ for $\ell=1–5$. Bottom panels: the corresponding Fourier expansion amplitudes plotted in the same two formats.
Figure 10: Heatmaps of the first four Fourier coefficients $C_{0}$$C_{3}$ for a Kerr black hole shadow, shown as functions of inclination $i$ and spin $a_{*}$ in the panels (top left to bottom right).
@@ IV.1 Forward Network: Shadow Generation from Parameters
It is important to note that we did not attempt to predict the Fourier coefficients $C_{\ell}$ directly from the physical parameters. Instead, the forward network is constructed to generate the full shadow profile $R(\psi)$, which mimics the type of data provided by observations. From these synthetic contours, the Fourier coefficients can then be extracted and used as inputs to the inverse network for parameter estimation. This two-step design ensures consistency with the observational pipeline, where images are first reconstructed and only afterward decomposed into invariant features.
@@ IV.2 Inverse Network: Parameter Estimation from Shadow Features
These degeneracies reflect a fundamental limitation of shadow-based parameter estimation rather than a shortcoming of the network itself. When the shadow morphology is nearly invariant under changes in the input parameters, no method based solely on the shadow contour can fully resolve the underlying spacetime properties. In such cases, the network naturally exhibits larger uncertainties, consistent with the intrinsic loss of information. A more complete inference framework would therefore need to combine shadow features with additional observables, such as polarization maps, variability signatures, or multi-wavelength constraints, in order to break these degeneracies.
@@ V Conclusion
Our analysis also highlights an important limitation: different parameter combinations can produce nearly indistinguishable shadows, leading to intrinsic degeneracies that no shadow-only method can resolve. The neural network captures this by showing larger uncertainties in such regions of parameter space. Breaking these degeneracies will require complementary observational information, for example from polarization measurements, variability studies, or independent mass and spin constraints.
@@ A.5 Kaluza–Klein Metric
$\frac{H_{1}}{M^{2}}$ $=\frac{8(b-2)(c-2)b}{(b+c)^{3}}+\frac{4(b-2)x}{b+c}+x^{2}$ (47)
$-\frac{2b\sqrt{(b^{2}-4)(c^{2}-4)}a_{*}\cos\theta}{(b+c)^{2}}+a_{*}^{2}\cos^{2}\theta,$
$\frac{H_{2}}{M^{2}}$ $=\frac{8(b-2)(c-2)c}{(b+c)^{3}}+\frac{4(c-2)x}{b+c}+x^{2}$
$-\frac{2c\sqrt{(b^{2}-4)(c^{2}-4)}a_{*}\cos\theta}{(b+c)^{2}}+a_{*}^{2}\cos^{2}\theta,$